OpenStudy (anonymous):

given matrix A= (1 -1 1 -1 2 -2 2 -2 1 -2 0 -1 4 2 1 0) prove that the determinant det A=0 without calculating the determinant.

OpenStudy (amistre64):

if one row or column is a scalar of another, I believe the det goes to 0

OpenStudy (anonymous):

R2=2R1

OpenStudy (amistre64):

R2 = R1*2

OpenStudy (amistre64):

yeah, what Gowt said :)

OpenStudy (anonymous):

do you think that this should be a sufficient answer? if R2= 2R1, det=0 (2 -2 2 -2)= 2( 1 -1 1 -1) = (2 -2 2 -2)

OpenStudy (amistre64):

hmmm, that seems more "teacher" dependant. If you have literature where you can cite the thrm used ... that might be a better route to take

OpenStudy (amistre64):

it has to do with linear depence/independence ....